OpenStudy (anonymous):

If f(x)=4sin(x)+3x^x, find f'(3).

hartnn (hartnn):

that x^x is a tricky part find its derivative separately using logarithmic differentiation. can you ?

OpenStudy (anonymous):

xln3x ?

hartnn (hartnn):

let me verify first is it \(3 (x^x)\) or \((3x)^x\) ?

OpenStudy (anonymous):

x(1/3x)*3?

OpenStudy (anonymous):

it doesnt specify in the problem its just \[3x ^{x}\]

hartnn (hartnn):

ok, so let y = x^x log y = log x^x = x log x try to differentiate log y = x log x using product rule on right and chain rule on left what do u get ?

OpenStudy (anonymous):

It should differerentiate to 3(ln(x)+1)x^x +4cos(x)

OpenStudy (anonymous):

@hartnn i get x(1/x) so 1

hartnn (hartnn):

no, you didn't use product rule on right ?

OpenStudy (anonymous):

@BigTeach2345 i got that an i plugged in 3 and he answer i get is wrong):

OpenStudy (anonymous):

@hartnn ln(x)+1

hartnn (hartnn):

correct! so 1/y y' = ln x +1 so, y' = y (1+ ln x) but y is x^x so, derivative of x^x is x^x (1+ln x) do you get this part ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Is your calculator in radians or degrees?

hartnn (hartnn):

so what about derivative of f(x) now ? f(x)=4sin(x)+3x^x f'(x) = ...?

OpenStudy (anonymous):

@gabyportillo95 I get 108.33 in radians and 116.284 in degrees

OpenStudy (anonymous):

its 3(ln(x)+1)x^x +4cos(x)

hartnn (hartnn):

thats correct, just plug in x=3 there

OpenStudy (anonymous):

i got it thank you! all i had to do was change it to radians :O

hartnn (hartnn):

i am glad you got it :)